Given that f(x)= 4x+4 and g(x)= x^2-3x-4, Find (gof)(4)
Answer by Aldorozos(172)

We replace x in g(x)= x^2-3x-4 with 4x+4
We get: (4x+4)^2 -3(4x+4) - 4
Now we replace x with 4
(4*4 +4)^2 -3(4*4 +4) -4 = 20^2 - 3(20)-4